. Explain how the margin of safety would change when per unit price increases by 5% but variable expense per unit decreases by more than 5%.
Answer #1

When per unit sales price is increased by 5% and variable expense per unit decrease by more than 5%, then margin of safety will increase. Because, sales price increase by 55 and variable expenses decrease by more than 5% will lead to increase in the contribution margin. When contribution margin is increased then sales required to break even will reduce, which will lead to increase in margin of safety. Let us understand this with the help of an example below:

Suppose, the following data for sale price, variable cost and fixed cost.

Sale price = $100

Variable cost = $50

Fixed cost = $10

Then

Contribution margin = Sale price - Variable cost = $100 - $50 = $50

Contribution margin ratio = Contribution margin / Sales * 100 = $50 / $100 * 100 = 50% or 0.5

Break even sales = Fixed cost / Contribution margin ratio = $10 / 0.5 = $20

Margin of safety = Sales - Break even sales

Margin of safety = $100 - $20 = $80

Now, suppose sale price increases by 5% and variable cost decreases by 10%, then new sale price and variable costs are:

New sale price = $105

New variable cost $40

Fixed cost = $10

New Contribution margin = New Sale price - New Variable cost = $105 - $40 = $65

New Contribution margin ratio = New Contribution margin / New Sales * 100 = $65 / $105 * 100 = 61.9% or 0.619

New Break even sales = Fixed cost / New Contribution margin ratio = $10 / 0.619 = $16.155

Margin of safety = New Sales - New Break even sales

Margin of safety = $105 - $16.155 = $88.845

So, from the above, we can see that earlier margin of safety was $80, but with increase in sale price and decrease in variable cost, margin of safety increases to $88.845